OpenStudy (anonymous):

answer the following and show your work X/X+3 +2/5 = 0

OpenStudy (phi):

is this \[ \frac{x}{x+3} + \frac{2}{5}=0 \]

OpenStudy (anonymous):

Try multiplying both sides by \(x+3\).